# Learning Neural Contracting Dynamics: Extended Linearization and Global Guarantees

### v3

@@ 1 Introduction
where $f(x)$ is an unknown (autonomous) vector field governing a dynamical process. Beyond approximating the vector field $f$, it is desirable to ensure that the learned vector field is well-behaved, e.g., it should admit a single stable equilibrium point and its behaviour should be robust in the face of uncertainty. Indeed, stability and robustness guarantees are crucial for many dynamical and control systems applications, including robotic manipulation, reinforcement learning, and forecasting.
@@
To ensure robustness in addition to stability, there has been increased interest in learning contracting dynamics. A dynamical system is said to be contracting if any two trajectories converge to one another exponentially quickly with respect to some metric [25]. It is known that if a system is contracting, it is also exponentially incrementally input-to-state stable [34]. Additionally, if the system is autonomous, it admits a unique equilibrium, is input-to-state stable, and has two Lyapunov functions that establish exponential stability of the equilibrium [9]. Since establishing contractivity globally requires satisfying a matrix partial differential inequality everywhere, prior works have focused on establishing contractivity in the neighborhood of training data.
@@ 1.1 Related Works
Stable, but not necessarily contracting dynamics. Numerous works have aimed to learn stable dynamical systems from data including [21, 27, 35, 29, 38, 7, 14]. In [21], the authors introduce the Stable Estimator of Dynamical Systems (SEDS), which leverages a Gaussian mixture model to approximate the dynamics and enforce asymptotic stability via constraints on learnable parameters. In [27], the authors jointly learn the dynamics and a Lyapunov function for the system and project the dynamics onto the set of dynamics which enforce an exponentially decay the Lyapunov function. This projection is done in closed-form and establishes global exponential convergence of the dynamics to the origin. In [35], the authors introduce Imitation Flow where trajectories are mapped to a latent space where states evolve in time according to a stable SDE. In [29], Euclideanizing Flows is introduced where the latent dynamics are enforced to follow natural gradient dynamics [2]. A similar approach is taken in [38] where additionally collision avoidance is considered.
@@
Existing works on learning contracting dynamics. Learning contracting vector fields from demonstrations has attracted attention due to robustness guarantees [30, 32, 33, 34, 5]. In [30], the dynamics are defined using a Gaussian mixture model and the contraction metric is parametrized to be a symmetric matrix of polynomial functions. Convergence to an equilibrium is established using partial contraction theory [36]. In [32], the dynamics are defined via an optimization problem over a reproducing kernel Hilbert space; the dynamics are constrainted to be locally contracting around the data points, i.e., there are points in state space where the dynamics are not contracting. In [33] and [34], the authors study controlled dynamical systems of the form $\dot{x}=f(x,u)$, where $u$ is a control input and train neural networks to find a feedback controller such that the closed-loop dynamics are approximately contracting, i.e., contractivity is not enforced but instead lack of contractivity is penalized in the cost function.
@@
Recent work, [5], has proposed a Neural Contractive Dynamical System (NCDS) which learns a dynamical system which is explicitly contracting to a trajectory. NCDS parameterizes contracting vector fields by learning symmetric Jacobians and performing line integrals to evaluate the underlying vector field. NCDS is computationally costly because of this integration. Additionally, the Jacobian parametrization used is overly restrictive, because not all contracting vector fields have symmetric Jacobians. Constraining the vector field to have symmetric Jacobian is equivalent to enforcing that the vector field is a negative gradient flow and contracting with respect to the identity metric [37]. For scalability to higher dimensions, NCDS leverages a latent space structure where an encoder maps the data space to a lower-dimensional space and enforces contracting dynamics in this latent space. Then a decoder “projects" the latent-space dynamics to the full data space. It is then argued that on the submanifold defined by the image of the decoder, the dynamics are contracting.
@@ 1.2 Contributions
In this paper, we present a novel model for learning deep dynamics with global contracting guarantees. We refer to this model as Extended Linearized Contracting Dynamics (ELCD). To the best of our knowledge, ELCD is the first model to ensure global contraction guarantees. To facilitate the development of this model, we provide a review of contracting dynamics and extended linearization of nonlinear dynamics. Leveraging extended linearization, we factorize our vector field as $f(x)=A(x,x^{*})(x-x^{*})$, where $x^{*}$ is the equilibrium of the dynamics. We enforce negative definiteness of the symmetric part of $A(x,x^{*})$ everywhere and prove (i) global exponential stability of $x^{*}$, (ii) equilibrium contractivity of our dynamics to $x^{*}$, and (iii) using a converse contraction theorem, contractivity of the dynamics with respect to some metric.
Since negative definiteness of the symmetric part of $A$ is not sufficient to capture all contracting dynamics, we introduce a latent space of equal dimension as the data space and learn diffeomorphisms between the data space and this latent space. The diffeomorphisms provide additional flexibility in the contraction metric and allow learning of more general contracting dynamics compared to those which are solely equilibrium contracting.
@@
Our model, ELCD, directly improves on NCDS in several ways. We parameterize the vector field directly instead of parameterizing its Jacobian. Doing so prevents us from needing to integrate the Jacobian to calculate the vector field and thus speeds up training and inference. ELCD is also more expressive than NCDS because it can represent vector fields with asymmetric Jacobians. Additionally, ELCD is guaranteed to be contracting to an arbitrary equilibrium point, either selected or learned, at all training steps. NCDS, in contrast, must learn the equilibrium point over the course of training. Additionally, NCDS learns an encoder and decoder for a lower-dimensional latent space and thus can only be contracting on the submanifold defined by the image of the decoder. From initial conditions that are not on this submanifold, NCDS may not exhibit contracting behavior. In contrast, since the latent space of ELCD is of the same dimension as the data space, we train diffeomorphisms that ensure global contractivity in the data space. ELCD exhibits better performance than NCDS at reproducing trajectories from the LASA [24], n-link Pendulum, and Rosenbrock datasets. We additionally compare against the Euclideanizing Flow [29], and Stable Deep Dynamics [27] models.
@@ 2 Preliminaries
The problem under consideration is as follows: given a demonstration $\mathcal{D}=\{x_{i},\dot{x}_{i}\}$ consisting of a set of $n$ state, $x_{i}\in{}^{d}$, and velocity, $\dot{x}_{i}\in{}^{d}$, pairs, we aim to learn a neural network $f(x_{i})=\dot{x}_{i}$ that parameterizes a globally contracting dynamical system with equilibrium point $x^{*}$, such that $f(x^{*})=0$. In essence, this task requires learning both the vector field and the contraction metric, $M$ such that they jointly satisfy (2).
@@ 3.1 Motivation
This factorization is referred to as extended linearization in analogy to standard linearization, i.e., for $x$ close to $x^{*}$, $f(x)\approx D\mkern-0.75mu{f}(x^{*})(x-x^{*})$. We remark that when $d\geq 2$, extended linearization is not unique, and for a given $f$, there may exist several $A$ such that $f(x)=A(x,x^{*})(x-x^{*})$. In other words, (8) showcases one valid choice for $A$ such that this factorization holds. Indeed, this nonuniqueness has been leveraged in some prior words, e.g. Section 3.1.3 in [34], to yield less conservative contractivity conditions.
@@
where $\frac{\partial A}{\partial x}:{}^{n}\times{}^{n}\rightarrow{}^{n\times n\times n}$ is a third-order tensor-valued mapping. In what follows, we will show that a more simple condition will imply contractivity of the vector field. Namely, negative definiteness of the symmetric part of $A(x,x^{*})$ will be sufficient for contractivity of the dynamical system $\dot{x}=A(x,x^{*})(x-x^{*})$.
@@ 3.2 Parametrization of $A(x,x^{*})$
We use the method of partial contraction as in [36]. To this end, define the virtual system
$\dot{y}=A(x,x^{*})(y-x^{*}).$ (13)
We will establish that this virtual system is contracting in the identity metric, i.e., (2) is satisfied with $M(x)=I_{d}$. We see that the Jacobian for this virtual system is simply $A(x,x^{*})$ and
$A(x,x^{*})+A(x,x^{*})^{\top}$ $=-2P_{s}(x,x^{*})^{\top}P_{s}(x,x^{*})-2\alpha I_{d}$
@@
In other words, in view of (3),and since $M(x)=I_{d}$, any two trajectories $y_{1}(t)$ and $y_{2}(t)$ of the virtual system satisfy
@@
Note that we can pick one trajectory to be $y_{1}(t)=x^{*}$ for all $t$ and we can pick $y_{2}(t)=x(t)$, where $x(t)$ is a trajectory of the nominal system. This then establishes the claim. ∎
@@
While it is challenging, in general, to compute the metric (15), numerical considerations for approximating it arbitrarily closely are presented in [16]. In practice, one would select $C(x)=I_{d}$ and evaluate all integrals numerically.
We remark that our parametrization for $A$ in (11) is similar to the parametrization for the Jacobian of $f$ that was presented in [5]. There are a few key differences. Notably, $A(x,x^{*})$ may be asymmetric while the Jacobian in equation (3) in [5] is always symmetric. Notably, since their Jacobian is symmetric and negative definite, the vector field $\dot{x}=f(x)$ is a negative gradient flow, $\dot{x}=-\nabla V(x)$, for some strongly convex $V$. On the contrary, our dynamics (10) can exhibit richer behaviors than negative gradient flows in view of the asymmetry in $A(x,x^{*})$. Additionally, since the dynamics in [5] with their parametrization can be represented as $\dot{x}=-\nabla V(x)$ for some strongly convex $V$, it is guaranteed to be contracting in the identity metric, $M(x)=I_{d}$ [37]. On the other hand, our dynamics (10) is not necessarily contracting in the identity metric and instead is contracting in a more complex metric given in (15).
@@ 3.3 Latent Space Representation
Latent space learning is possible because contraction is invariant under differential coordinate changes. From Theorem 2 of [26], given a dynamical system $\dot{x}=f(x)$, $f:{}^{d}\to{}^{d}$, with $f$ satisfying (1), the system will also be contracting under the coordinate change $z=\phi(x)$ if $\phi:{}^{d}\rightarrow{}^{d}$ is a smooth diffeomorphism. Specifically, if $\dot{x}=f(x)$ is contracting with metric $M$, then the system that evolves $z$ is contracting as well with metric given by $\tilde{M}(x)=D\mkern-0.75mu{\phi}(x)^{-\top}M(x)D\mkern-0.75mu{\phi}(x)^{-1}$, where $z=\phi(x)$ and $D\mkern-0.75mu{\phi}$ is the Jacobian of the coordinate transform. In other words, We can learn vector fields that are contracting in an arbitrary metric by training a vector field $f$ which is parametrized as (10) and using a coordinate transform $\phi$.
NCDS, [5], treats the coordinate transform as a Variational Autoencoder (VAE) [22]. Their training procedure consists of two steps: first training the coordinate transform with VAE training, then training the function $f$ in the new learned coordinates.
@@ 3.5 Choice of Diffeomorphism
There are several popular neural network diffeopmorphisms including coupling layers [11, 29], normalizing flows [35], $\mathcal{M}$-flow [8, 5], spline flows [13], and radial basis functions [6].
@@ 3.6 Training
Our task is to simultaneously learn the contracting system $f(x)$ and the metric in which $f$ contracts. As previously discussed, the metric is implicitly determined by $\phi(x)$. [5] uses a two-step method. First, they treat $\phi$ as a variational autoencoder and minimize the evidence lower bound (ELBO). They then fix the encoder and train the model to evolve the state in latent space. Note that the VAE objective is to make the encoded data to match a standard-normal distribution. Notably, VAE training does not encourage the encoder to transform the data to space in which the data is corresponds to trajectories of a contracting system. If the data is not contracting in the transformed space, a contracting model will not be able to fully fit the data. This explains why, in our implementation of the two step-training, we are unable to reasonably learn the data. For further comparison with NCDS, we will train the encoder and model jointly.
@@ 4 Experiments
For all datasets we compare our method against NCDS  [5], Stable Deep Dynamics (SDD) [27] and Euclideanizing Flow (Eflow) [29]. See A.2 for a detailed discussion of the methods. We report the dynamic time warping distance (DTWD) [19] (see A.1) and standard deviation between predicted and data trajectories. See A.4 for model implementation details. See also the Appendix for more details on these models.
@@ 4.1 Datasets
Table 1: DTWD on LASA, Pendulum, and Rosenbrock datasets
@@
LASA-2D 0.37 $\pm$ 0.32 1.05 $\pm$ 0.25 0.59 $\pm$ 0.61 0.46 $\pm$0.57
LASA-4D 2.49 $\pm$ 2.4 2.24 $\pm$ 0.12 2.19 $\pm$ 1.23 1.34 $\pm$ 0.95
LASA-8D 5.26 $\pm$ 0.50 2.66 $\pm$ 0.63 5.04 $\pm$ 0.77 2.41 $\pm$ 0.68
Pendulum-4D 0.49 $\pm$ 0.11 0.17 $\pm$ 0.01 1.35 $\pm$ 2.260.16 $\pm$ 0.05
Pendulum-8D 0.75 $\pm$ 0.08 0.33 $\pm$ 0.01 2.88 $\pm$ 0.69 0.28 $\pm$ 0.10
Pendulum-16D 1.86 $\pm$ 0.14 0.45 $\pm$ 0.01 1.65 $\pm$ 0.31 0.51 $\pm$ 0.13
Rosenbrock-8D NaN 1.90 $\pm$ 0.16 2.74 $\pm$ 0.15 1.24 $\pm$ 0.26
Rosenbrock-16D NaN 3.57 $\pm$ 0.66 3.68 $\pm$ 0.12 2.26 $\pm$ 0.51

### v4

@@ 1 Introduction
where $f(x)$ is an unknown (autonomous) vector field governing a dynamical process. Beyond approximating the vector field $f$, it is desirable to ensure that the learned vector field is well-behaved. In many robotic tasks like grasping and navigation, a well-behaved system should always reach a fixed endpoint. Ideally, a learned system will still stably reach the desired endpoint even when pushed away from demonstration trajectories. Additionally, the learned system should robustly reach the desired endpoint in the face of uncertainty. In tasks such as manufacturing, animation, and human-robot interaction, functionality and safety require the learned system must smoothly follow a specific trajectory to its target.
@@
To ensure robustness and to allow for smooth trajectory following, there has been increased interest in learning contracting dynamics from data [6]. A dynamical system is said to be contracting if any two trajectories converge to one another exponentially quickly with respect to some metric [27]. If a learned contracting system that admits a demonstration trajectory is pushed off that trajectory, it will follow a new trajectory that exponentially converges to the demonstration trajectory. Additionally, if a system is contracting, it is exponentially incrementally input-to-state stable [39]. If the system is autonomous, it also admits a unique equilibrium, is input-to-state stable, and has two Lyapunov functions that establish exponential stability of the equilibrium [9]. Since establishing contractivity globally requires satisfying a matrix partial differential inequality everywhere, prior works have focused on establishing contractivity in the neighborhood of training data [35, 36].
@@ 1.1 Related Works
Stable, but not necessarily contracting dynamics. Numerous works have aimed to learn stable dynamical systems from data including [22, 30, 40, 32, 45, 7, 15]. In [22], the authors introduce the Stable Estimator of Dynamical Systems (SEDS), which leverages a Gaussian mixture model to approximate the dynamics and enforce asymptotic stability via constraints on learnable parameters. In [30], the authors jointly learn the dynamics and a Lyapunov function for the system and project the dynamics onto the set of dynamics which enforce an exponentially decay of the Lyapunov function. This projection is done in closed-form and establishes global exponential convergence of the dynamics to the origin. In [40], the authors introduce Imitation Flow where trajectories are mapped to a latent space where states evolve in time according to a stable SDE. In [32], Euclideanizing Flows is introduced where the latent dynamics are enforced to follow natural gradient dynamics [2]. A similar approach is taken in [45] where additionally collision avoidance is considered.
@@
Existing works on learning contracting dynamics. Learning contracting vector fields from demonstrations has attracted attention due to robustness guarantees [33, 35, 37, 39, 6]. In [33], the dynamics are defined using a Gaussian mixture model and the contraction metric is parametrized to be a symmetric matrix of polynomial functions. Convergence to an equilibrium is established using partial contraction theory [41]. In [35], the dynamics are defined via an optimization problem over a reproducing kernel Hilbert space; the dynamics are constrained to be locally contracting around the data points, i.e., there may be points in state space where the dynamics are not contracting. In [37] and [39], the authors study controlled dynamical systems of the form $\dot{x}=f(x,u)$, where $u$ is a control input and train neural networks to find a feedback controller such that the closed-loop dynamics are approximately contracting, i.e., contractivity is not enforced but instead lack of contractivity is penalized in the cost function.
@@
Recent work, [6], has proposed a Neural Contractive Dynamical System (NCDS) which learns a dynamical system which is explicitly contracting to a trajectory. NCDS parametrizes contracting vector fields by learning symmetric Jacobians and performing line integrals to evaluate the underlying vector field. NCDS is computationally costly because of this integration. Additionally, the Jacobian parametrization used is overly restrictive, because not all contracting vector fields have symmetric Jacobians. Constraining the vector field to have symmetric Jacobian is equivalent to enforcing that the vector field is a negative gradient flow and contracting with respect to the identity metric [42]. For scalability to higher dimensions, NCDS leverages a latent space structure where an encoder maps the data space to a lower-dimensional space and enforces contracting dynamics in this latent space. Then a decoder “projects" the latent-space dynamics to the full data space. It is then argued that on the submanifold defined by the image of the decoder, the dynamics are contracting.
@@ 1.2 Contributions
In this paper, we present a novel model for learning deep dynamics with global contraction guarantees. We refer to this model as Extended Linearized Contracting Dynamics (ELCD). To the best of our knowledge, ELCD is the first model to ensure global contraction guarantees. To facilitate the development of this model, we provide a review of contracting dynamics and extended linearization of nonlinear dynamics. Leveraging extended linearization, we factorize our vector field as $f(x)=A(x,x^{*})(x-x^{*})$, where $x^{*}$ is the equilibrium of the dynamics. We enforce negative definiteness of the symmetric part of $A(x,x^{*})$ everywhere and prove (i) global exponential stability of $x^{*}$, (ii) equilibrium contractivity of our dynamics to $x^{*}$, and (iii) using a converse contraction theorem, contractivity of the dynamics with respect to some metric.
Since negative definiteness of the symmetric part of $A$ is not sufficient to capture all contracting dynamics, we introduce a latent space of equal dimension as the data space and learn diffeomorphisms between the data space and this latent space. The diffeomorphisms provide additional flexibility in the contraction metric and allow learning of arbitrary contracting dynamics compared to those which are solely equilibrium contracting.
@@
Our model, ELCD, directly improves on NCDS in several ways. We parameterize the vector field directly instead of parametrizing its Jacobian. Doing so prevents us from needing to integrate the Jacobian to calculate the vector field and thus speeds up training and inference. ELCD is also more expressive than NCDS because it can represent vector fields with asymmetric Jacobians. Additionally, ELCD is guaranteed to be contracting to an arbitrary equilibrium point, either selected or learned, at all training steps. NCDS, in contrast, must learn the equilibrium point over the course of training. Additionally, NCDS learns an encoder and decoder for a lower-dimensional latent space and thus can only be contracting on the submanifold defined by the image of the decoder. From initial conditions that are not on this submanifold, NCDS may not exhibit contracting behavior. In contrast, since the latent space of ELCD is of the same dimension as the data space, we train diffeomorphisms that ensure global contractivity in the data space. ELCD exhibits better performance than NCDS at reproducing trajectories from the LASA [26], n-link Pendulum, and Rosenbrock datasets. We additionally compare against the Euclideanizing Flow [32], and Stable Deep Dynamics [30] models.
@@ 2 Preliminaries
The problem under consideration is as follows: given a set of demonstrations $\mathcal{D}=\{(x_{i},\dot{x}_{i})\}_{i=1}^{n}$ consisting of a set of $n$ state, $x_{i}\in{}^{d}$, and velocity, $\dot{x}_{i}\in{}^{d}$, pairs, we aim to learn a neural network $f(x_{i})=\dot{x}_{i}$ that parametrizes a globally contracting dynamical system with equilibrium point $x^{*}$, such that $f(x^{*})=0$. In essence, this task requires learning both the vector field and the contraction metric, $M$ such that they jointly satisfy (2).
@@ 3.1 Motivation
This factorization is referred to as extended linearization in analogy to standard linearization, i.e., for $x$ close to $x^{*}$, $f(x)\approx D\mkern-0.75mu{f}(x^{*})(x-x^{*})$. We remark that when $d\geq 2$, extended linearization is not unique, and for a given $f$, there may exist several $A$ such that $f(x)=A(x,x^{*})(x-x^{*})$. In other words, (8) showcases one valid choice for $A$ such that this factorization holds. Indeed, this nonuniqueness has been leveraged in some prior works, e.g. Section 3.1.3 in [39], to yield less conservative contractivity conditions.
@@
where $\frac{\partial A}{\partial x}:{}^{d}\times{}^{d}\rightarrow{}^{d\times d\times d}$ is a third-order tensor-valued mapping. In what follows, we will show that a more simple condition will imply contractivity of the vector field. Namely, negative definiteness of the symmetric part of $A(x,x^{*})$ will be sufficient for contractivity of the dynamical system $\dot{x}=A(x,x^{*})(x-x^{*})$.
@@ 3.2 Parametrization of $A(x,x^{*})$
We use the method of partial contraction as in [41]. Let $x(t)$ be a solution to $\dot{x}(t)=A(x(t),x^{\star})(x(t)-x^{*})$ with initial condition $x(0)=x_{0}$. Then, define the time-varying virtual system
$\dot{y}(t)=A(x(t),x^{*})(y(t)-x^{*}).$ (13)
We will establish that this virtual system is contracting in the identity metric, i.e., (2) is satisfied with $M(x)=I_{d}$. We see that the Jacobian for this virtual system is simply $A(x(t),x^{*})$ and
$A(x(t),x^{*})+A(x(t),x^{*})^{\top}$ $=-2P_{s}(x(t),x^{*})^{\top}P_{s}(x(t),x^{*})-2\alpha I_{d}$
@@
In other words, in view of (3), and since $M(x)=I_{d}$, any two solution trajectories $y_{1}(t)$ and $y_{2}(t)$ of the virtual system satisfy
@@
Note that we can pick one trajectory to be $y_{1}(t)=x^{*}$ for all $t$ and we can pick $y_{2}(t)=x(t)$. Since $x_{0}$ was arbitrary, this argument establishes the claim. ∎
@@
While it is challenging, in general, to compute the metric (15), numerical considerations for approximating it arbitrarily closely are presented in [17]. In practice, one would select $C(x)=I_{d}$ and evaluate all integrals numerically. For ELCD, unless one directly needs to know the contraction metric for application purposes, it is not required to compute the contraction metric at any point during either training or inference.
We remark that our parametrization for $A$ in (11) is similar to the parametrization for the Jacobian of $f$ that was presented in [6]. There are however a few key differences. Notably, $A(x,x^{*})$ may be asymmetric while the Jacobian in equation (3) in [6] is always symmetric. Notably, since the Jacobian in [6] is symmetric and negative definite, the vector field $\dot{x}=f(x)$ is a negative gradient flow, $\dot{x}=-\nabla V(x)$, for some strongly convex function $V$. On the contrary, our dynamics (10) can exhibit richer behaviors than negative gradient flows in view of the asymmetry in $A(x,x^{*})$. Additionally, since the dynamics in [6] with their parametrization can be represented as $\dot{x}=-\nabla V(x)$ for some strongly convex $V$, it is guaranteed to be contracting in the identity metric, $M(x)=I_{d}$ [42]. On the other hand, our dynamics (10) is not necessarily contracting in the identity metric and instead is contracting in a more complex metric given in (15).
@@ 3.3 Latent Space Representation
Latent space learning is possible because contraction is invariant under differential coordinate changes. From Theorem 2 of [29], given a dynamical system $\dot{x}=f(x)$, $f:{}^{d}\to{}^{d}$, with $f$ satisfying (1), the system will also be contracting under the coordinate change $z=\phi(x)$ if $\phi:{}^{d}\rightarrow{}^{d}$ is a smooth diffeomorphism. Specifically, if $\dot{x}=f(x)$ is contracting with metric $M$, then the system that evolves $z$ is contracting as well with metric given by $\tilde{M}(z)=D\mkern-0.75mu{\phi}(z)^{-\top}M(z)D\mkern-0.75mu{\phi}(z)^{-1}$, where $z=\phi(x)$ and $D\mkern-0.75mu{\phi}$ is the Jacobian of the coordinate transform. In other words, We can learn vector fields that are contracting in an arbitrary metric by training a vector field $f$ which is parametrized as (10) and using a coordinate transform $\phi$.
NCDS [6], treats the coordinate transform as a Variational Autoencoder (VAE) [23]. Their training procedure consists of two steps: first training the coordinate transform with VAE training, then training the function $f$ in the new learned coordinates.
@@ 3.5 Choice of Diffeomorphism
There are several popular neural network diffeomorphisms including coupling layers [12, 32], normalizing flows [40], $\mathcal{M}$-flow [8, 6], spline flows [14], and radial basis functions [5].
@@ 3.6 Training
Our task is to simultaneously learn the contracting system $f(x)$ and the metric in which $f$ contracts. As previously discussed, the metric is implicitly determined by $\phi(x)$. [6] uses a two-step method. First, they treat $\phi$ as a variational autoencoder and maximize the evidence lower bound (ELBO). They then fix the encoder and train the model to evolve the state in latent space. Note that the VAE objective is to make the encoded data to match a standard-normal distribution. Notably, VAE training does not encourage the encoder to transform the data to space in which the data corresponds to trajectories of a contracting system. If the data is not contracting in the transformed space, a contracting model will not be able to fully fit the data. This explains why, in our implementation of the two step-training, we are unable to reasonably learn the data. For further comparison with NCDS, we will train the encoder and model jointly.
@@ 4 Experiments
The code for our model and data can be found here: https://github.com/seanjaffe1/Extended-Linearized-Contracting-Dynamics. For all datasets we compare our method against NCDS  [6], Stable Deep Dynamics (SDD) [30] and Euclideanizing Flow (EFlow) [32]. See A.2 for a detailed discussion of the methods. We report the dynamic time warping distance (DTWD) [20] (see A.1) and standard deviation between predicted and data trajectories. See A.5 for model implementation details. See also the Appendix for more details on these models.
@@ 4.1 Datasets
Table 1: Mean DTWD $\pm$ one standard deviation across 10 runs on LASA, multi-link pendulum, and Rosenbrock datasets
@@
LASA-2D 0.37 $\pm$ 0.32 1.05 $\pm$ 0.25 0.59 $\pm$ 0.61 0.12 $\pm$0.11
LASA-4D 2.49 $\pm$ 2.4 2.24 $\pm$ 0.12 2.19 $\pm$ 1.23 0.80 $\pm$ 0.54
LASA-8D 5.26 $\pm$ 0.50 2.66 $\pm$ 0.63 5.04 $\pm$ 0.77 1.52 $\pm$ 0.61
Pendulum-4D 0.49 $\pm$ 0.11 0.17 $\pm$ 0.01 1.35 $\pm$ 2.260.03 $\pm$ 0.01
Pendulum-8D 0.75 $\pm$ 0.08 0.33 $\pm$ 0.01 2.88 $\pm$ 0.69 0.14 $\pm$ 0.03
Pendulum-16D 1.86 $\pm$ 0.14 0.45 $\pm$ 0.01 1.65 $\pm$ 0.31 0.44 $\pm$ 0.09
Rosenbrock-8D NaN 1.90 $\pm$ 0.16 2.74 $\pm$ 0.15 1.22 $\pm$ 0.01
Rosenbrock-16D NaN 3.57 $\pm$ 0.66 3.68 $\pm$ 0.12 2.57 $\pm$ 0.09